The measures of the two acute angles of a right triangle differ by 18 degrees. What are their measures?
Answer by RAY100(1637)

Let x and y be angle measures
.
x+y=90,,,,,given rt angle
.
x-y=18,,,,,given
.
adding relationships
.
2x =108
.
x=54,,,,,,,,,,,,,y= 90-54 = 36
.
.
check,,,,,54-36=18,,,,,ok